OpenStudy (anonymous):

The bumper of a certain model of car is approved to withstand a collision speed of up to 3 meters per second. How many miles per hour is this equivalent to?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (gorv):

1 meter = 0.00062 miles 3 meter = 0.00062*3=???

OpenStudy (gorv):

3600 sec = 1hour 1 sec= 1/3600 hour

OpenStudy (gorv):

replace meter and sec with this data